OpenStudy (anonymous):

simplify sqrt{128}+sqrt{1682}*sqrt{196}-sqrt{1152}...please help me

OpenStudy (tyteen4a03):

sqrt(196) = 14. First prime factorize the numbers then factorize the whole expression.

OpenStudy (anonymous):

@experimentX help??

OpenStudy (experimentx):

try factorizing ... this should be your answer http://www.wolframalpha.com/input/?i=sqrt {128}%2Bsqrt{1682}*sqrt{196}-sqrt{1152}

OpenStudy (anonymous):

i am unable to factorize :(

OpenStudy (experimentx):

128 = 2*64 = 2 * 2*32 = 2*2*2*16 = 2*2*2*2*2*2*2

OpenStudy (anonymous):

@experimentX thanks...i guess i can do the rest now

OpenStudy (experimentx):

yep ... try factoring for others too.